# Mixture of Experts Meets Prompt-Based Continual Learning

### v1

@@ Appendix B Proof of Theoretical Results
Appendix C Training Algorithm of HiDe-Prompt
@@
1:Pre-trained transformer backbone $f_{\theta}$, training sets $\mathcal{D}_{1},...,\mathcal{D}_{T}$ , number of tasks $T$, number of epochs $E$, hyper-parameters $\alpha$, $\tau$ and $\lambda$.
2:Parameters $\bm{p}_{1},...,\bm{p}_{T},\omega$ and $\psi$
@@
17:         Obtain $\mathcal{G}_{c}$ from $f_{\theta},\bm{p}_{t}$ and $\mathcal{D}_{t}$triangleright Instructed Representations      return $(\bm{p}_{1},...,\bm{p}_{T},\omega,\psi)$
@@
HiDe-Prompt maintains an expandable pool of task-specific prompts $\bm{e}_{t}$, each optimized for a specific task $\mathcal{D}_{t}$ using a cross-entropy loss within the WTP objective. To prevent forgetting, previous prompts $\bm{e}_{1},..,\bm{e}_{t-1}$ remain frozen. Knowledge transfer across tasks is facilitated by a prompt ensemble (PE) strategy: the current prompt is initialized with the last one $\bm{e}_{t}\leftarrow\bm{e}_{t-1}$ and refined using a weighted combination of all past prompts $\bm{p}_{t}=\alpha\sum_{i=1}^{t-1}\bm{e}_{i}+(1-\alpha)\bm{e}_{t}$, where $\alpha$ is a hyper-parameter. Notably, HiDe-Prompt incorporates contrastive regularization within the WTP objective, pushing features of the new task away from those of past tasks represented by the prototypes of old class distributions $\mathcal{G}_{c}$. Let $\mathcal{H}_{t}=\{f_{\theta}(\bm{x}_{i}^{(t)},\bm{p}_{t})|\;i=1,..,N_{t}\}$ be the embedding transformation of $\mathcal{D}_{t}$ and $\bm{\mu}_{c}$ be the mean of $\mathcal{G}_{c}$. The contrastive loss can be written as
@@
where $\mathcal{H}_{i,c}$ is constructed by sampling an equal number of pseudo representations from $\mathcal{G}_{c}$ for $c\in\mathcal{Y}^{(i)}$ and $i=1,...,t$.
@@
where ^$\hat{\mathcal{H}}_{c}$ is constructed by sampling an equal number of pseudo representations from $\hat{\mathcal{G}}_{c}$ for $c\in\mathcal{Y}^{(i)}$ and $i=1,...,t$. Please refer to Algorithm 1 for more details.
Appendix D Experimental Details
@@
Datasets We use commonly-used datasets in the field of continual learning, including (1) Split CIFAR-100 [18]: This dataset comprises images from 100 classes. These classes are divided randomly into 10 separate incremental tasks, with each task featuring a distinct set of classes. (2) Split ImageNet-R [18]: This dataset is composed of images from 200 classes. It includes challenging examples from the original ImageNet [33] dataset and newly gathered images representing diverse styles. These classes are also randomly divided into 10 distinct incremental tasks. (3) Split CUB-200 [39]: This dataset consists of fine-grained images of 200 different bird species. It is randomly divided into 10 incremental tasks, each comprising a unique class subset. (4) 5-Datasets [9]: This composite dataset incorporates CIFAR-10 [18], MNIST [19], Fashion-MNIST [46], SVHN [26], and notMNIST [3]. Each of these datasets is treated as a separate incremental task, permitting for the assessment of the effects of significant variations between tasks.
@@
Prompt-Based Approaches We compare NoRGa against recent prompt-based continual learning approaches: L2P [45], DualPrompt [44], CODA-Prompt [35], S-Prompt [43] and HiDe-Prompt [40]. To ensure a fair comparison, we replicate these methods using the configurations reported in their respective papers. S-Prompt in the original paper trains a separate prompt and classifier head for each task. At evaluation, it infers domain id as the nearest centroid obtained by applying K-Means on the training data. To adapt S-Prompt to CIL, we use one common classifier head for all tasks. For NoRGa, we adopt the same configuration as HiDe-Prompt, which utilizes Prefix Tuning [21] as its prompt-based methodology. Learnable scalar factors $\alpha$ and $\tau$ are frozen after the first task’s training to mitigate catastrophic forgetting. We further optimize NoRGa by selecting the best non-linear activation function $\sigma$ via cross-validation among $\tanh$, $\mathrm{sigmoid}$, and $\mathrm{GELU}$.
@@
Evaluation Metric We employ three common metrics to measure the performance the methods, including final average accuracy (FA), cumulative average accuracy (CA), and average forgetting measure (FM). Let $S_{i,t}$ denote the accuracy on the $i$-th task after learning the $t$-th task, and $A_{t}$ represent the average accuracy as $A_{t}=\frac{1}{t}\sum_{i=1}^{t}S_{i,t}$. Upon learning all $T$ tasks, we compute $\mathrm{FA}=A_{T}$, $\mathrm{CA}=\frac{1}{T}\sum_{t=1}^{T}A_{t}$, and $\mathrm{FM}=\frac{1}{T-1}\sum_{i=1}^{T-1}\mathrm{max}_{t\in\{1,\dots,T-1\}}(S_{i,t}-S_{i,T})$. It’s worth noting that FA and CA are prioritized over FM, as they inherently encompass both plasticity and forgetting, with FM providing supplementary context [35].
@@
Implementation Details We train and test on one NVIDIA A100 GPU for baselines and our method. We leverage a pre-trained ViT-B/16 model as the backbone. Training employs an Adam optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$), a batch size of 128, and a constant learning rate of 0.005 for all methods except CODA-Prompt. CODA-Prompt utilizes a cosine decaying learning rate starting at 0.001. Additionally, a grid search technique was implemented to determine the most appropriate number of epochs for effective training.
Appendix E Task-incremental learning results
@@
Because HiDe-Prompt optimizes prompt parameters specifically for within-task prediction (WTP), NoRGa inherently aligns with this objective, leading to generally better continual learning performance. We demonstrate this improvement through experiments in a task-incremental learning setting, where task labels are available during inference (as in Table 4). While HiDe-Prompt performs well, NoRGa shows consistent improvement across all scenarios. Notably, NoRGa with sigmoid activation achieves the highest final average accuracy in both Split CIFAR-100 and Split CUB-200 with Sup-21K training. Additionally, NoRGa demonstrates its effectiveness even with self-supervised pretraining, further solidifying its advantage over the original Prefix Tuning model. Overall, NoRGa variants outperform HiDe-Prompt on both datasets and under both training conditions.

### v2

@@ Appendix B Proof of Theoretical Results
Appendix C Discussion of related Mixture of Experts works
Recently, the MoE model has been employed to mitigate catastrophic forgetting in continual learning. For example, [58] focused on continual learning in vision-language models by adapting a pre-trained vision-language model to new tasks through learning a mixture of specialized adapter modules. [58] introduced an MoE structure onto a frozen CLIP, utilizing a mixture of adapters to modify the MLP block after the MSA layer. In contrast, our work centers on general continual learning with pre-trained models, leveraging the inherent MoE architecture of MSA layers. Consequently, our MoE model placement differs from that of [58]. By employing prefix tuning, we demonstrate that it is analogous to introducing new prefix experts to scale and adapt these pre-trained MoE models to downstream tasks. Furthermore, while [58] utilizes task-specific routers, our approach employs task-specific prompts that encapsulate both task-specific router and expert parameters.
The parameters cost is usually considered in practical memory-constrained continual learning scenarios. Dynamic routing mechanism can be employed for gating-based neural networks [18]. To improve the parameter efficiency of the final model, we can integrate this mechanism in the proposed method. Specifically, each head in the MSA layers comprises $N$ MoE models, where $N$ is the length of the input sequence. This allows for a dynamic routing mechanism to enhance parameter efficiency. For instance, [18] proposed a dynamic routing strategy that adaptively adjusts the number of activated experts based on the input. The computation for any MoE model’s gating is directly correlated with the corresponding row in the attention matrix, which encapsulates the MoE model’s score functions. For example, selecting the top $k$ experts via Top-K routing in the $i$-th MoE model is equivalent to identifying the top $k$ largest values in the $i$-th row of the attention matrix. To implement [18], we first sort the elements in the $i$-th row from highest to lowest, then find the smallest set of experts whose cumulative probability exceeds the threshold. Consequently, unselected experts remain inactive, reducing the need to compute all elements of the value matrix within self-attention.
Appendix D Training Algorithm of HiDe-Prompt
@@
1:Pre-trained transformer backbone $f_{\theta}$, training sets $\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ , number of tasks $T$, number of epochs $E$, hyper-parameters $\alpha$, $\tau$ and $\lambda$.
2:Parameters $\bm{p}_{1},\dots,\bm{p}_{T},\omega$ and $\psi$
@@
17:         Obtain $\mathcal{G}_{c}$ from $f_{\theta},\bm{p}_{t}$ and $\mathcal{D}_{t}$triangleright Instructed Representations      return $(\bm{p}_{1},\dots,\bm{p}_{T},\omega,\psi)$
@@
HiDe-Prompt maintains an expandable pool of task-specific prompts $\bm{e}_{t}$, each optimized for a specific task $\mathcal{D}_{t}$ using a cross-entropy loss within the WTP objective. To prevent forgetting, previous prompts $\bm{e}_{1},\dots,\bm{e}_{t-1}$ remain frozen. Knowledge transfer across tasks is facilitated by a prompt ensemble (PE) strategy: the current prompt is initialized with the last one $\bm{e}_{t}\leftarrow\bm{e}_{t-1}$ and refined using a weighted combination of all past prompts $\bm{p}_{t}=\alpha\sum_{i=1}^{t-1}\bm{e}_{i}+(1-\alpha)\bm{e}_{t}$, where $\alpha$ is a hyper-parameter. Notably, HiDe-Prompt incorporates contrastive regularization within the WTP objective, pushing features of the new task away from those of past tasks represented by the prototypes of old class distributions $\mathcal{G}_{c}$. Let $\mathcal{H}_{t}=\{f_{\theta}(\bm{x}_{i}^{(t)},\bm{p}_{t})|\;i=1,\dots,N_{t}\}$ be the embedding transformation of $\mathcal{D}_{t}$ and $\bm{\mu}_{c}$ be the mean of $\mathcal{G}_{c}$. The contrastive loss can be written as
@@
where $\mathcal{H}_{i,c}$ is constructed by sampling an equal number of pseudo representations from $\mathcal{G}_{c}$ for $c\in\mathcal{Y}^{(i)}$ and $i=1,\dots,t$.
@@
where ^$\hat{\mathcal{H}}_{c}$ is constructed by sampling an equal number of pseudo representations from $\hat{\mathcal{G}}_{c}$ for $c\in\mathcal{Y}^{(i)}$ and $i=1,\dots,t$. Please refer to Algorithm 1 for more details.
Appendix E Experimental Details
@@
Datasets. We use commonly-used datasets in the field of continual learning, including (1) Split CIFAR-100 [23]: This dataset comprises images from 100 classes. These classes are divided randomly into 10 separate incremental tasks, with each task featuring a distinct set of classes. (2) Split ImageNet-R [23]: This dataset is composed of images from 200 classes. It includes challenging examples from the original ImageNet [40] dataset and newly gathered images representing diverse styles. These classes are also randomly divided into 10 distinct incremental tasks. (3) Split CUB-200 [48]: This dataset consists of fine-grained images of 200 different bird species. It is randomly divided into 10 incremental tasks, each comprising a unique class subset. (4) 5-Datasets [9]: This composite dataset incorporates CIFAR-10 [23], MNIST [24], Fashion-MNIST [55], SVHN [31], and notMNIST [3]. Each of these datasets is treated as a separate incremental task, permitting for the assessment of the effects of significant variations between tasks.
@@
Prompt-Based Approaches. We compare NoRGa against recent prompt-based continual learning approaches: L2P [54], DualPrompt [53], CODA-Prompt [42], S-Prompt [52] and HiDe-Prompt [49]. To ensure a fair comparison, we replicate these methods using the configurations reported in their respective papers. S-Prompt in the original paper trains a separate prompt and classifier head for each task. At evaluation, it infers domain id as the nearest centroid obtained by applying K-Means on the training data. To adapt S-Prompt to CIL, we use one common classifier head for all tasks. For NoRGa, we adopt the same configuration as HiDe-Prompt, which utilizes Prefix Tuning [26] as its prompt-based methodology. Learnable scalar factors $\alpha$ and $\tau$ are frozen after the first task’s training to mitigate catastrophic forgetting. We further optimize NoRGa by selecting the best non-linear activation function $\sigma$ via cross-validation among $\tanh$, $\mathrm{sigmoid}$, and $\mathrm{GELU}$.
@@
Evaluation Metric. We employ three common metrics to measure the performance the methods, including final average accuracy (FA), cumulative average accuracy (CA), and average forgetting measure (FM). Let $S_{i,t}$ denote the accuracy on the $i$-th task after learning the $t$-th task, and $A_{t}$ represent the average accuracy as $A_{t}=\frac{1}{t}\sum_{i=1}^{t}S_{i,t}$. Upon learning all $T$ tasks, we compute $\mathrm{FA}=A_{T}$, $\mathrm{CA}=\frac{1}{T}\sum_{t=1}^{T}A_{t}$, and $\mathrm{FM}=\frac{1}{T-1}\sum_{i=1}^{T-1}\mathrm{max}_{t\in\{1,\dots,T-1\}}(S_{i,t}-S_{i,T})$. It’s worth noting that FA and CA are prioritized over FM, as they inherently encompass both plasticity and forgetting, with FM providing supplementary context [42].
@@
Implementation Details. We train and test on one NVIDIA A100 GPU for baselines and our method. We leverage a pre-trained ViT-B/16 model as the backbone. Training employs an Adam optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$), a batch size of 128, and a constant learning rate of 0.005 for all methods except CODA-Prompt. CODA-Prompt utilizes a cosine decaying learning rate starting at 0.001. Additionally, a grid search technique was implemented to determine the most appropriate number of epochs for effective training.
Appendix F Task-incremental Learning Results
@@
Because HiDe-Prompt optimizes prompt parameters specifically for within-task prediction (WTP), NoRGa inherently aligns with this objective, leading to generally better continual learning performance. We demonstrate this improvement through experiments in a task-incremental learning setting, where task labels are available during inference (as in Table 4). While HiDe-Prompt performs well, NoRGa shows consistent improvement across all scenarios. Notably, NoRGa with sigmoid activation achieves the highest final average accuracy in both Split CIFAR-100 and Split CUB-200 with Sup-21K training. Additionally, NoRGa demonstrates its effectiveness even with self-supervised pre-training, further solidifying its advantage over the original prefix tuning model. Overall, NoRGa variants outperform HiDe-Prompt on both datasets and under both training conditions.
Appendix G Comparison to Different Parameter-Efficient Fine-Tuning Methods
Table 5: Performance comparison of different PEFT methods using ViT-B/16 with Sup-21K weights. Here we present Final Average Accuracy (FA).
Method Split CIFAR-100 Split CUB-200
HiDe-Prompt 92.61 86.56
HiDe-LoRA 92.71 87.37
HiDe-Adapter 92.73 87.10
NoRGa 94.48 90.90
As the advantages of different parameter-efficient fine-tuning (PEFT) methods remain an open question, we briefly describe them through our revealed connection between self-attention and MoE. Prefix tuning introduces additional parameters at the input of MSA layers to adapt the pre-trained model representation, contrasting with Adapter [16], which insert adaptive parameters between layers, often replacing MLP blocks. LoRA [17] approximates weight updates with low-rank matrices and adds them to the backbone weights. Our work shows that the MSA layer in a pre-trained model can be seen as a pre-trained MoE architecture. Applying LoRA to the MSA layer refines both the pre-trained experts and their corresponding score functions for downstream tasks. In contrast, prefix tuning expands the pre-trained MoE models by incorporating new experts while preserving the original components, rather than modifying the pre-trained experts like LoRA.
NoRGa emerges as a simple, parameter-efficient fine-tuning method and can be regarded as a distinct implementation of prompts. However, our novel perspective on the interplay between self-attention, prefix tuning, and mixture of experts enables us to theoretically substantiate the effectiveness of NoRGa as shown in Section 4.
For empirical comparison, we integrate the framework of HiDe-Prompt with different PEFT techniques and Sup-21K weights, evaluating performance on Split CIFAR-100 and Split CUB-200. The results are summarized in Table 5. The table shows that NoRGa consistently outperforms the other PEFT methods on both datasets, suggesting its effectiveness. Nevertheless, further investigation with LoRA and Adapter would be necessary to draw more definitive conclusions.
While exploring alternative PEFT methods might offer improvements in WTP performance, these approaches lack theoretical guarantees and could lead to an increased number of parameters. In contrast, our NoRGa method modifies the original score functions of prefix tuning to enhance WTP performance with theoretical rigor. Importantly, NoRGa maintains the same parameter count as HiDe-Prompt, which is crucial in CL due to memory constraints.
Appendix H Comparison with Pre-trained Model-based Methods
Table 6: Performance comparison of pre-trained model-based continual learning methods using ViT-B/16 with Sup-21K weights. Here we present Final Average Accuracy (FA).
Method Split CIFAR-100 Split CUB-200
ADAM + VPT-D 85.04 85.28
ADAM + SSF 85.27 85.67
ADAM + Adapter 87.29 85.84
RanPAC 92.20 90.30
NoRGa 94.48 90.90
Table 7: Ablation study on the effect of learnable $\alpha$ and $\tau$ with Sup-21K weights. Here we present Final Average Accuracy (FA).
Method Split CIFAR-100 Split CUB-200
HiDe-Prompt 92.61 86.56
Learnable $\alpha$, Fixed $\tau$ 94.38 90.45
Fixed $\alpha$, Learnable $\tau$ 94.42 90.48
Fixed $\alpha$, Fixed $\tau$ 94.29 90.32
Learnable $\alpha$, Learnable $\tau$ 94.48 90.90
Previous works have demonstrated that utilizing pre-trained models (PTM) significantly enhances performance for continual learning, often surpassing the performance of non-PTM-based methods. Moreover, studies have shown that first-task adaptation and simple PEFT-style tuning can achieve competitive performance [20, 37, 60, 29] with prompt-based methods. For instance, [20] demonstrated that appending a nearest class mean (NCM) classifier to a ViT model’s feature outputs, can serve as a strong baseline. [37, 60] enhanced this strategy by adapting the pre-trained model to the first task using the three PEFT methods for transformer networks [60] and the FiLM method for CNNs [37]. Additionally, [29] improved NCM by incorporating second-order statistics—covariance and Gram matrices. However, these methods, which fine-tune only the backbone for the initial task, may not always ensure satisfactory separation of new tasks’ features. Our work focuses on continually adapting the backbone, utilizing task-specific prompts to consistently capture emerging tasks’ characteristics, and proposing a novel method to enhance the CL performance of previous prompting methods.
To validate our approach, we compare it against state-of-the-art PTM-based continual learning methods, including ADAM [60] and RanPAC [29], using Split CIFAR-100 and Split CUB-200 datasets. The results are summarized in Table 6. In comparison to other PTM-based continual learning methods, NoRGa demonstrates competitive performance across both evaluated datasets. For instance, on Split CIFAR-100, NoRGa achieves an FA of 94.48%, exceeding the next best method by over 2%. Similarly, on Split CUB-200, NoRGa delivers strong results relative to other baselines. These improvements highlight the effectiveness of our method in mitigating catastrophic forgetting and preserving knowledge retention across multiple tasks.
Appendix I Efficiency Tests
Figure 3: Validation loss on Split CUB-200 throughout the training of the first task.
We compare the validation loss of NoRGa and HiDe-Prompt throughout the first task on Split CUB-200, as illustrated in Figure 3. The results demonstrate that NoRGa consistently outperforms HiDe-Prompt throughout the training process. This empirical evidence supports the theoretical advantages of NoRGa over HiDe-Prompt.
Appendix J Effect of Learnable Hyperparameters
As described in Section 4.1 and Appendix E, in our framework, $\alpha$ and $\tau$ are learnable hyperparameters and optimized through backpropagation during the first task, eliminating the need for manual tuning. Additionally, our theoretical analysis of NoRGa’s statistical efficiency in Section 4.2 holds for any values of $\alpha$ and $\tau$, demonstrating the theoretical robustness. To further investigate, we evaluated both fixed and learnable settings for these hyperparameters. For the fixed case, we set their values to 1. We report FA on Split CUB-200 and Split CIFAR-100 with Sup-21K weights. The results are summarized in Table 7. Although performance slightly decreased with fixed hyperparameters, it still outperforms HiDe-Prompt, indicating our method’s empirical robustness.
Appendix K Training Times
Table 8: Comparison of training times for HiDe-Prompt and NoRGa. All experiments were conducted on a single NVIDIA A100 GPU.
Method Split CIFAR-100 Split ImageNet-R Split CUB-200 5-Datasets
HiDe-Prompt 2.80h 2.67h 1.04h 24.06h
NoRGa 2.85h 2.70h 1.10h 24.23h
We utilize a single A100 GPU for all experiments. The training times are summarized in Table 8. While NoRGa exhibits slightly longer training times compared to HiDe-Prompt, it consistently achieves significantly better performance. This demonstrates the effectiveness of NoRGa while maintaining competitive training efficiency.